# Dynamic Attention-Guided Diffusion for Image Super-Resolution

### v3

@@ Abstract
Diffusion models in image Super-Resolution (SR) treat all image regions with uniform intensity, which risks compromising the overall image quality. To address this, we introduce “You Only Diffuse Areas” (YODA), a dynamic attention-guided diffusion method for image SR. YODA selectively focuses on spatial regions using attention maps derived from the low-resolution image and the current time step in the diffusion process. This time-dependent targeting enables a more efficient conversion to high-resolution outputs by focusing on areas that benefit the most from the iterative refinement process, i.e., detail-rich objects. We empirically validate YODA by extending leading diffusion-based methods SR3 and SRDiff. Our experiments demonstrate new state-of-the-art performance in face and general SR across PSNR, SSIM, and LPIPS metrics. A notable finding is YODA’s stabilization effect by reducing color shifts, especially when training with small batch sizes.
@@ 1 Introduction
The goal of image Super-Resolution (SR) is to enhance Low-Resolution (LR) into High-Resolution (HR) images [35]. This process has a significant impact in fields such as medical imaging, remote sensing, and consumer electronics [33, 16, 44]. Despite its long history, image SR remains a fascinating yet challenging domain due to its inherently ill-posed nature: any given LR image can lead to several valid HR images, and vice versa [2, 40]. Thanks to deep learning, image SR has made significant progress [13]. Initial regression-based methods, such as early convolutional neural networks, work great at low magnification ratios[27, 7, 43]. However, they fail to produce high-frequency details at high magnification ratios and generate over-smoothed images [34].
@@
Recently, diffusion models have emerged with better human-rated quality compared to regression-based methods, but they also introduce new challenges [21, 38, 46, 9]. Their indiscriminate processing of image regions leads to computational redundancies and suboptimal enhancements. Some recent methods have reduced computational demands by working in latent space like LDMs [37], by exploiting the relationship between LR and HR latent representations like PartDiff [48], or by starting with a better-initialized forward diffusion instead of pure Gaussian noise like in CCDF [10]. However, there is still a lack of strategies that adapt the model capacity based on the spatial importance of image regions.
In this paper, we question the common approach of diffusion models for image SR: Is it necessary to update the entire image at every time step in the reverse diffusion process? We hypothesize that not all image regions require the same level of detail enhancement. For instance, a face in the foreground may need more attention than a simple, monochromatic background. Recognizing this variability in the need for detail enhancement underscores a critical inefficiency in traditional diffusion methods. Treating all image regions with uniform intensity risks compromising the overall image quality. To bridge this gap, we introduce “You Only Diffuse Areas” (YODA), an efficient diffusion mechanism focusing on detail-rich areas using time-dependent and attention-guided masking. YODA starts by obtaining an attention map, highlighting regions that need more refinement. After identification, YODA systematically replaces the highlighted regions with SR predictions during the denoising process, depending on the current time step. As a result, detail-rich areas are refined more often. Our approach is analogous to inpainting methods like RePaint [31], where only a pre-defined masked region is updated to generate complementing content. In YODA’s case, however, the selection of regions to update is time-dependent. We design a dynamic approach that creates expanding masks, starting from detail-rich regions and converging towards enhancing the overall image. A key advantage of YODA is its compatibility with existing diffusion models, allowing for a plug&play application. We integrate YODA with SR3 [38] for face SR and SRDiff [25] for general SR, achieving notable improvements in image quality. Interestingly, YODA also improves the training process. For instance, our face SR experiments require a significantly reduced batch size due to limited hardware access. When training with smaller batch sizes, SR3 also suffers from color shifts [42, 8], while YODA produces color distributions faithfully. In summary, our work:
@@
  • introduces YODA, an attention-guided diffusion approach that emphasizes important image areas through masked refinement. As a result, it refines detail-rich areas more often, which leads to higher overall image quality.
@@
  • demonstrates that attention-guided diffusion results in better training conditions, accurate color predictions, and better perceptual quality.
@@
  • shows that YODA improves the training performance when using smaller batch sizes, which is crucial in limited hardware scenarios.
@@ 2 Background
Our method uses attention maps for attention-guided diffusion. Thus, this section introduces the main components: DDPMs [21] and the DINO framework [6].
@@ 2.1 DDPMs
Denoising Diffusion Probabilistic Models (DDPMs) employ two distinct Markov chains [21]: the first models the forward diffusion process $q$ transitioning from an input $\mathbf{x}$ to a pre-defined prior distribution with intermediate states $\mathbf{z}_{t}$, $0<t\leq T$, while the second models the backward diffusion process $p$, reverting from the prior distribution back to the intended target distribution $p\left(\mathbf{z}_{0}\mid\mathbf{z}_{T},\mathbf{x}\right)$. In image SR, we designate $\mathbf{x}$ as the LR image and the target $\mathbf{z}_{0}$ as the desired HR image. The prior distribution is generally set manually, e.g., Gaussian noise.
2.1.1 Forward Diffusion
In the forward diffusion phase, an HR image $\mathbf{z}_{0}$ is incrementally modified by adding Gaussian noise over a series of time steps. This process can be mathematically represented as:
@@ 2.1.1 Forward Diffusion
where $\gamma_{t}=\prod_{i=1}^{t}(1-\alpha_{i})$ The intermediate step $\mathbf{z}_{t}$ is derived by:
@@
2.1.2 Backward Diffusion
The backward diffusion process is where the model learns to denoise, effectively reversing the forward diffusion to recover the HR image. In image SR, the reverse process is conditioned on the LR image to guide the generation of the HR image:
@@ 2.1.2 Backward Diffusion
2.1.3 Optimization
The optimization goal for DDPMs is to train the parameterized model to accurately predict the noise added during the diffusion process. The loss function used to measure the accuracy of the noise prediction is:
@@ 2.2 DINO
DINO is a self-supervised learning approach for feature extractors on unlabeled data [6]. It employs a teacher and a student network, where the student learns to imitate the features learned by the teacher. The student gets only local views of the image (i.e., $96\times 96$), whereas the teacher receives global views (i.e., $224\times 224$). This setup encourages the student to learn “local-to-global” correspondences. The features learned through self-supervision are directly accessible in the self-attention modules. These self-attention maps provide information on the scene layout and important object boundaries. We leverage the generality, availability, and robustness of these attention maps as a measure of an image’s saliency to guide the diffusion process. In another context, a similar approach has been applied to image compression, demonstrating its ability to capture essential image content in the attention maps [3]. More details can be found in the supplementary material.
@@ 3 Methodology
This section describes the components of our proposed method, “You Only Diffuse Areas” (YODA). YODA is divided into three main phases:
  • Identifying Key Regions: Estimate the importance of pixel positions in a LR image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ with an attention map $\mathbf{A}\in\mathbb{R}^{H\times W}$.
  • Time-Dependent Masking: Using $\mathbf{A}$, define time-dependent, binary masks $\mathbf{M}:\mathbb{N}_{0}\to\{0,1\}^{H\times W}$ to identify salient areas at time step $t\in\mathbb{N}_{0}$ of the diffusion process.
  • Guided Backward Diffusion: Condition the diffusion process on the regions identified by the time-dependent masking $\mathbf{M}(t)$.
@@ 3.1 Identifying Key Regions
Super-resolving with YODA begins by identifying areas of an image with greater details. This is achieved by generating an attention map $\mathbf{A}$ with $0\leq\mathbf{A}_{i,j}\leq 1$ from the LR image $\mathbf{x}$. The greater the value of $\mathbf{A}_{i,j}$, the more important it is. Note that $\mathbf{A}$ has to be generated only once for each image.
For generating $\mathbf{A}$, we evaluated several approaches, including innate methods (i.e., not-learnable) and learnable methods, i.e., ResNet and Transformer architectures [20, 14]. For the latter, we leverage the DINO framework for its robustness in self-supervised learning, extracting refined attention maps directly from LR images without necessitating extra annotated data [6]. This choice is motivated by DINO’s demonstrated efficacy in highlighting essential features within images using pre-existing models, e.g., for image compression [3]. An additional overview of how DINO and the attention maps are used is shown in the supplementary materials. Next, we describe the process of creating time-dependent masks for backward diffusion, utilizing the attention map $\mathbf{A}$.
@@ 3.2 Time-Dependent Masking
Given the LR input image $\mathbf{x}$ and the attention map $\mathbf{A}$, we introduce a strategy to focus the diffusion process to detail-rich areas. Each position in $\mathbf{A}$ indicates its semantic importance, influencing the number of refinement steps it receives. Thus, for two positions $(i,j)$ and $(i^{\prime},j^{\prime})$ with $\mathbf{A}_{i,j}>\mathbf{A}_{i^{\prime},j^{\prime}}$, YODA applies more refinement steps to the location $(i,j)$ than to $(i^{\prime},j^{\prime})$. Since $0\leq\mathbf{A}_{i,j}\leq 1$, the number of diffusion steps employed to a specific position $(i,j)$ is determined as a proportion of the maximum time steps, $T$. For instance, $\mathbf{A}_{i,j}=0.7$ means $(i,j)$ is refined during 70% of all diffusion steps. In addition, a lower bound hyperparameter $10<l<$ 1 ensures that every region undergoes a minimum level of refinement. In other words, the hyperparameter $l$ guarantees that every spatial position is refined at least $l\cdot T$ times. Note that the backward diffusion process starts from time step $T$ and ends in time step 0. We define the time-dependent masking process at time step $0\leq t\leq T$ with $t\to 0$ as follows:
@@
This equation ensures that the diffusion process gets applied a variable number of times for different regions, allowing the salient areas to diffuse over a longer time span. It is important to highlight that once a spatial position is marked for refinement, it continues to undergo refinement across all subsequent steps as $t$ approaches 0: $\mathbf{M}(t)_{i,j}\geq\mathbf{M}(t-k)_{i,j}\,\forall k>0$. Figure 1 shows an example of our time-dependent masking. For each time step $t$, we can determine with $\mathbf{M}(t)_{i,j}=1$ whether a given spatial position $(i,j)$ should be refined or not.
Figure 1: Overview of YODA. First, extract an attention map $\mathbf{A}$ from the LR input. Next, use the values of $\mathbf{A}$ to produce a time-dependent masking $\mathbf{M}(t)$. For $t:T\to 0$, the area of selected pixels expands from detail-rich regions to the whole image. Our diffusion process uses these masks for dynamic and attention-guided refinement, emphasizing important regions. More specifically, it starts with masked areas that need refinement (derived from $\mathbf{z}_{t}$ and $\mathbf{M}(t)$) and LR regions, which retain the noise level needed for the next time step. Finally, the SR and LR areas are combined to form a whole image with no masked-out regions for the next iteration.
@@ 3.3 Guided Backward Diffusion
YODA’s guided diffusion process iteratively refines the image from a noisy state $\mathbf{z}_{T}$ to a high-resolution state $\mathbf{z}_{0}$. This phase involves selectively refining areas based on the current time step’s mask, $\mathbf{M}(t)$, and blending these refined areas with the unrefined, remaining LR regions. YODA ensures a seamless transition between refined and unrefined areas, improving image quality with a focus on key regions.
At each time step $t$, the areas that will be refined for transitioning from $t$ to $(t-1)$ are determined based on the current iteration $\mathbf{z}_{t}$ and the current mask $\mathbf{M}(t)$:
@@
Next, we divide the image into two components that will later be combined as the output for the next time step: $\mathbf{z}^{SR}_{t-1}$, which is the refined image prediction, and $\mathbf{z}^{LR}_{t-1}$, the complementary LR image. The state $\mathbf{z}^{LR}_{t-1}$ represents unchanged LR areas but is sampled using $\mathbf{x}$ as the mean. Both components acquire the same noise level $\Sigma_{\theta}(\mathbf{\widetilde{z}}_{t},\mathbf{x},\gamma_{t})$, and can be described by:
@@
Finally, YODA combines the complementing, non-overlapping image regions [^1] This formulation is similar to RePaint [31], a method for diffusion-based inpainting. While RePaint uses a constant mask for all time steps, YODA uses a time-dependent procedure that enables dynamic control over the refined image regions.:
@@ 3.4 Optimization
To confine the backward diffusion process to specific image regions as determined by the current time step $0\leq t\leq T$ and the corresponding mask $\mathbf{M}(t)$, YODA improves the training objective Equation 7 as follows:
@@
The loss function focuses on the highlighted spatial regions within the mask $\mathbf{M}(t)$. Consequently, YODA optimizes only areas described by $\mathbf{M}(t)$.

### v4

@@ Abstract
Diffusion models in image Super-Resolution (SR) treat all image regions uniformly, which risks compromising the overall image quality by potentially introducing artifacts during denoising of less-complex regions. To address this, we propose “You Only Diffuse Areas” (YODA), a dynamic attention-guided diffusion process for image SR. YODA selectively focuses on spatial regions defined by attention maps derived from the low-resolution images and the current denoising time step. This time-dependent targeting enables a more efficient conversion to high-resolution outputs by focusing on areas that benefit the most from the iterative refinement process, i.e., detail-rich objects. We empirically validate YODA by extending leading diffusion-based methods SR3, DiffBIR, and SRDiff. Our experiments demonstrate new state-of-the-art performances in face and general SR tasks across PSNR, SSIM, and LPIPS metrics. As a side effect, we find that YODA reduces color shift issues and stabilizes training with small batches.
@@ 1 Introduction
The goal of image Super-Resolution (SR) is to enhance Low-Resolution (LR) into High-Resolution (HR) images [41]. Improvements to this field significantly impact many applications, like medical imaging, remote sensing, and consumer electronics [38, 18, 52]. Despite its long history, image SR remains a fascinating yet challenging domain due to its inherently ill-posed nature: any LR image can lead to several valid HR images, and vice versa [2, 48]. Thanks to deep learning, SR has made significant progress [15]. Initial regression-based methods, such as early convolutional neural networks, work great at low magnification ratios [30, 8, 54]. However, they fail to produce high-frequency details at high magnification ratios ($\geq$ 4) and generate over-smoothed results [40]. Such scale ratios require models capable of hallucinating realistic details that fit the overall image.
@@
Recently, generative diffusion models have emerged with better human-rated quality compared to regression-based methods, but they also introduced new challenges [23, 46, 56, 10]. Their indiscriminate processing of image regions leads to computational redundancies and suboptimal enhancements. Some recent methods address the first issue and reduce computational demands by working in latent space like LDMs [45], by exploiting the relationship between LR and HR latents like PartDiff [62], or by starting with a better-initialized forward diffusion instead of pure noise like in CCDF [11]. Yet, strategies to adapt model capacity based on spatial importance remain underexplored.
This paper takes the first step toward addressing the second issue and challenges the common approach of SR diffusion models by asking: Do we need to update the entire image at every time step? We hypothesize that not all image regions require the same level of detail enhancement. For instance, a face in the foreground may need more refinement than a simple, monochromatic background. Recognizing this variability in the need for detail enhancement underscores a critical inefficiency in traditional diffusion methods. Treating all image regions uniformly risks compromising the overall image quality by introducing artifacts and shifts to low-complex regions. Unlike methods that target computational efficiency, we aim to boost image quality by minimizing distortions across different low-complex regions.
In response, we introduce a diffusion mechanism focusing on detail-rich areas using time-dependent and attention-guided masking. Our method, coined ”You Only Diffuse Areas” (YODA), starts by obtaining an attention map that highlights regions that need more refinement. After identification, YODA systematically replaces highlighted regions with SR predictions during the denoising process. In particular, regions with high attention values (detail-rich & salient) are refined more often. Our approach is analogous to inpainting methods like RePaint [35], where only a pre-defined region is updated to generate complementing content. In YODA’s case, however, the selected regions are time-dependent. To that end, we design a dynamic approach that creates expanding masks, starting from detail-rich regions and converging toward the overall image.
A key advantage of YODA is its compatibility with existing diffusion models, allowing for a plug&play application. We integrate YODA with three models: SR3 [46] and DiffBIR [31] for face SR and SRDiff [28] for general SR. Interestingly, YODA achieves notable image quality improvements and also improves the training process. When training with smaller batch sizes, SR3 suffers from color shifts [50, 9] while YODA produces faithful color distributions. In summary, our work:
@@
  • introduces YODA, an attention-guided diffusion approach that emphasizes image areas through masked refinement. Thus, it refines detail-rich areas more often, which leads to higher image quality.
@@
  • demonstrates that attention-guided diffusion results in better training conditions, accurate color reproduction, and competitive perceptual quality results.
@@
  • reveals that YODA improves the training performance when using smaller batch sizes, which is crucial in limited hardware scenarios.
@@ 2 Background
Our method uses attention maps for attention-guided diffusion. We leverage the self-supervised DINO framework [7] to extract attention maps. Thus, this section introduces the main components: DDPMs [23] and the DINO [7]. We refer to the supplementary materials for a discussion of related methods, such as other diffusion approaches and spatial-selection SR methods.
@@ 2.1 DDPMs
Denoising Diffusion Probabilistic Models (DDPMs) employ two distinct Markov chains [23]: the first models the forward diffusion process $q$ transitioning from an input $\mathbf{x}$ to a pre-defined prior distribution with intermediate states $\mathbf{z}_{t}$, $0<t\leq T$, while the second models the backward diffusion process $p$, reverting from the prior distribution back to the intended target distribution $p\left(\mathbf{z}_{0}\mid\mathbf{z}_{T},\mathbf{x}\right)$. In image SR, we designate $\mathbf{x}$ as the LR image and the target $\mathbf{z}_{0}$ as the desired HR image. The prior distribution is usually Gaussian noise.
Forward Diffusion: In forward diffusion, an HR image $\mathbf{z}_{0}$ is incrementally modified by adding Gaussian noise over a series of time steps. This process can be mathematically represented as:
@@
where $\gamma_{t}=\prod_{i=1}^{t}(1-\alpha_{i})$ The intermediate step $\mathbf{z}_{t}$ is
@@
Backward Diffusion: The backward diffusion process is where the model learns to denoise, effectively reversing the forward diffusion to recover the HR image. In image SR, the reverse process is conditioned on the LR image to guide the generation of the HR image:
@@
Optimization: The optimization goal for DDPMs is to train the parameterized model to accurately predict the noise added during the diffusion process. The loss function used to measure the accuracy of the noise prediction is:
@@ 2.2 DINO
DINO is a self-supervised learning approach for feature extractors on unlabeled data [7]. It employs a teacher and a student network, where the student learns to imitate the features learned by the teacher. The student gets only local views of the image (i.e., $96\times 96$), whereas the teacher receives global views (i.e., $224\times 224$). This setup encourages the student to learn “local-to-global” correspondences. The features learned through self-supervision are directly accessible in the self-attention modules. These self-attention maps provide information on the scene layout and object boundaries. We leverage the generality, availability, and robustness of these attention maps as a measure of an image’s saliency to guide the diffusion process for significantly improved image quality. In another context, a similar approach has been applied to image compression, demonstrating its ability to capture essential image content in the attention maps [3].
Figure 1: Overview of YODA. First, extract an attention map $\mathbf{A}$ from the LR input. Next, use the values of $\mathbf{A}$ to produce a time-dependent masking $\mathbf{M}(t)$. For $t:T\to 0$, the area of selected pixels expands from detail-rich regions to the whole image. Our diffusion process uses these masks for dynamic and attention-guided refinement, emphasizing regions differently. More specifically, it starts with masked areas that need refinement (derived from $\mathbf{z}_{t}$ and $\mathbf{M}(t)$) and LR regions, which retain the noise level needed for the next time step. Finally, the SR and LR areas are combined to form a whole image with no masked-out regions for the next iteration.
@@ 3 Methodology
Our proposed method, coined “You Only Diffuse Areas” (YODA), has three major phases:
  • Identifying Key Regions: Estimate the weighting of pixel positions in a LR image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ with an attention map $\mathbf{A}\in\mathbb{R}^{H\times W}$.
  • Time-Dependent Masking: Use $\mathbf{A}$ to define a time-dependent, binary mask generator function $\mathbf{M}:\mathbb{N}_{0}\to\{0,1\}^{H\times W}$. The generated masks identify salient areas at time step $t\in\mathbb{N}_{0}$ of the diffusion process.
  • Guided Backward Diffusion: Concentrate the diffusion process on the regions identified by the time-dependent masking $\mathbf{M}(t)$ and generate a partially enhanced image by combining the prediction with complementing LR areas.
@@ 3.1 Identifying Key Regions
YODA starts by prioritizing areas in the input. This is achieved by generating an attention map $\mathbf{A}$ with $0\leq\mathbf{A}_{i,j}\leq 1$ from the LR image $\mathbf{x}$. The greater the value of $\mathbf{A}_{i,j}$, the more refinements it receives. Note that extracting $\mathbf{A}$ is computationally efficient as it has to be generated only once for each image. For generating $\mathbf{A}$, we evaluated several approaches, including innate methods (i.e., not-learnable) and learnable methods, i.e., ResNet [22] and Transformer architectures [16]. For the latter, we leverage the DINO framework for its robustness in self-supervised learning, extracting refined attention maps directly from LR images without necessitating extra annotated data [7]. This choice is motivated by DINO’s demonstrated efficacy in highlighting essential features within images using pre-existing models, e.g., for image compression [3].
Note that it is challenging to define important regions in SR because there is no clear definition. However, we observe that foreground objects are typically critical to human perception, while background areas seem less significant. Therefore, we choose to consider self-supervised methods to extract attention maps, which serve as an unbiased proxy for identifying objects and weightings in an image. This choice is based on the observation that self-supervised methods like DINO naturally focus on regions that generally capture human attention [7]. Our experiments with YODA confirm that emphasizing these areas improves performance, as later results will demonstrate. An additional overview of how DINO and the attention maps are used is shown in the supplementary materials. Next, we describe the process of creating time-dependently masks for the backward diffusion by utilizing the attention map $\mathbf{A}$.
@@ 3.2 Time-Dependent Masking
Given the LR input image $\mathbf{x}$ and the attention map $\mathbf{A}$, we introduce a novel strategy to dynamically focus the diffusion process on salient areas. Even though $\mathbf{A}$ is fixed, we will use it to refocus the diffusion model during the backward diffusion process dynamically. Thus, we can leverage it to influence the number of refinement steps for each position. Therefore, for two positions $(i,j)$ and $(i^{\prime},j^{\prime})$ with $\mathbf{A}_{i,j}>\mathbf{A}_{i^{\prime},j^{\prime}}$, YODA applies more refinement steps to the location $(i,j)$ than to $(i^{\prime},j^{\prime})$. Since $0\leq\mathbf{A}_{i,j}\leq 1$, the number of diffusion steps employed to a specific position $(i,j)$ is determined as a proportion of the maximum time steps, $T$. For instance, $\mathbf{A}_{i,j}=0.7$ means $(i,j)$ is refined during 70% of all diffusion steps. In addition, we introduce a lower bound hyperparameter $10<l<$ 1, ensuring that every region undergoes a minimum amount of refinements. In other words, the hyperparameter $l$ reliably guarantees that every spatial position is refined at least $l\cdot T$ times. As the backward diffusion process progresses from time step $T$ to 0, we can define the time-dependent masking process for any time step $T\geq t\geq 0$ approaching $t=0$ as:
@@
Equation 8 ensures that the diffusion process gets applied a variable number of times for different regions, allowing the salient areas to diffuse over a longer time span. It is important to highlight that once a spatial position is marked for refinement, it continues to undergo refinement across all subsequent steps: $\mathbf{M}(t)_{i,j}\geq\mathbf{M}(t-k)_{i,j}\,\forall k>0$. Figure 1 shows an example of our time-dependent masking. For each time step $t$, we can determine with $\mathbf{M}(t)_{i,j}=1$ whether a given spatial position $(i,j)$ should be refined or not.
@@ 3.3 Guided Backward Diffusion
YODA’s guided diffusion process iteratively refines the image from a noisy state $\mathbf{z}_{T}$ to a HR state $\mathbf{z}_{0}$. This phase involves selectively refining areas based on the current time step’s mask, $\mathbf{M}(t)$, and blending these refined areas with the unrefined, remaining LR regions, $(1-\mathbf{M}(t))$. YODA ensures a seamless transition between refined and unrefined areas, improving image quality with a focus on key regions.
More specifically, at each time step $t$, the areas that will be refined when transitioning from $t$ to $(t-1)$are determined based on the current iteration $\mathbf{z}_{t}$ and the current mask $\mathbf{M}(t)$:
@@
[…]
@@
Finally, YODA combines the complementing and non-overlapping image regions into a full image [^1] Equation 12 is similar to RePaint [35], a diffusion-based inpainting method While RePaint uses a constant mask for all time steps, YODA has time-dependent masks to dynamically control the updated image regions.:
@@
Note that we use $\mathbf{\widetilde{z}}_{t}$ instead of $\mathbf{z}_{t}$ in Equation 10 and Equation 11. In our initial experiments, masking before the noise prediction (i.e., $\mathbf{\widetilde{z}}_{t}$) produced a marginal improvement in comparison to the full intermediate state $\mathbf{z}_{t}$ (around 0.1-0.2 dB in PSNR). We theorize that it is connected to the optimization target explained next. With $\mathbf{\widetilde{z}}_{t}$, we force the model inherently to focus locally, which, due to our selective loss function, would otherwise have to be learned.
@@ 3.4 Optimization
To confine the backward diffusion process to specific image regions as determined by the current time step $0\leq t\leq T$ and the corresponding mask $\mathbf{M}(t)$, we adapt the training objective from Equation 7 as follows to focus on regions within the mask $\mathbf{M}(t)$. Thus, YODA optimizes only areas described by $\mathbf{M}(t)$: